# On Viscosity Solutions of Hamilton-Jacobi Equations in the Wasserstein space and the Vanishing Viscosity Limit

### v1

@@ 4. Useful Estimates and Lemmata
We first start by showing that, as a consequence of the Benamou-Brenier control representation, the elements of the subdifferential of $W_{2}$-Lipschitz functions are bounded:
@@
($L^{2}$ version of [DS, Lemma  3.4]) Let $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a $W_{2}$- Lipschitz function and set $L=\mathrm{Lip}(U;W_{2})<\infty$. Let $\mu_{0}\in\mathcal{P}_{2,ac}(\mathbb{R}^{d})$ be such that $\mathcal{E}(\mu_{0})<\infty$ and let $\delta\geq 0$, and $p\in\partial^{+}(U-\delta\mathcal{E})(\mu_{0})$ (resp. $p\in\partial^{-}(U+\delta\mathcal{E})(\mu_{0})$), then $\|p+\delta\nabla\log\mu_{0}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\leq L$ (resp. $\|p-\delta\nabla\log\mu_{0}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\leq L$)
@@
Since the proof is very similar to the one given in [DS, Lemma  3.4], we only give a sketch of it in the superdifferential case.
@@
(4.1) $\partial_{t}\mu_{t}+\operatorname{div}(v(x)\mu_{t})=0\quad\text{in }(t_{0},\infty),\quad\mu|_{t=t_{0}}=\mu_{0}.$
@@
Moreover, by the representation theorem for solutions of the continuity equation given in [AGS, Thm. 8.3.1] we have
@@
Moreover, by the renormalization property satisfied by $v$ and standard regularization and localization procedures (see [DS, Lemma  3.4]), we also have
@@
Fix now $h,\varepsilon>0$ and $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$. By similar estimates performed in [DS, Lemma  3.4], it follows, from the superdifferential property and the Cauchy-Schwarz inequality, that for all sufficiently small $h>0$, depending on $\|v\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}$ and $\varepsilon$,
@@
is bounded in $L^{2}_{\mu_{0}}(\mathbb{R}^{d})$ with operator norm $L$. By the $L_{\mu_{0}}^{2}(\mathbb{R}^{d})$ duality, this element belongs to $L^{2}_{\mu_{0}}(\mathbb{R}^{d})$ and $\|p+\delta\nabla\log\mu_{0}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\leq L$. ∎
The following Lemma will be directly applied in the proof of the vanishing viscosity theorem.
Lemma 4.2 (Analogue of Proposition 3.1 of [DS] in our context.).
Let $U,V:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be two bounded functions s.t. $\mathrm{Lip}(U;W_{1})+\mathrm{Lip}(V;W_{2})<\infty$ and suppose $V$ is l.s.c. with respect to the $W_{p}$ topology for some $p\in[1,2)$. Fix $\alpha,\delta>0$, then the function $\Phi:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})\to[-\infty,+\infty)$ defined by
$\Phi(\mu,\nu)=U(\mu)-V(\nu)-\frac{1}{2\alpha}W_{2}(\mu,\nu)-\delta\mathcal{E}(\mu)-\delta\pi\mathcal{M}_{2}(\mu)-\delta\pi\mathcal{M}_{2}(\nu)$
@@
    where $\sigma\in\Gamma_{0}(\overline{\nu},\overline{\mu})$ is such that $(\pi_{2},\pi_{1})_{\#}\sigma=(\mathrm{Id}\times T_{\overline{\mu}}^{\overline{\nu}})_{\#}\overline{\mu}\in\Gamma_{0}(\overline{\mu},\overline{\nu}).$
    Then $p\in{\partial^{+}}\big(U-\delta\mathcal{E}\big)(\overline{\mu})$ and $\xi\in\partial^{-}V(\overline{\nu})$. Moreover,
    $\|p+\delta\nabla\log\overline{\mu}\|_{L^{\infty}}\leq\mathrm{Lip}(U;W_{1}),\,\|p+\delta\nabla\log\overline{\mu}\|_{L^{2}_{\overline{\mu}}(\mathbb{R}^{d})}\leq\mathrm{Lip}(U;W_{2})\text{ and }\|\xi\|_{\overline{\nu}}\leq\mathrm{Lip}(V;W_{2}).$
@@
The attainment of the maximum follows the same argument used in the proof of [DS, Proposition 3.1]. The moment penalization forces any maximizing sequence to remain in a compact (for the weak convergence) region of $\mathcal{P}_{2}(\mathbb{R}^{d})$, so that it admits a weakly convergent subsequence. By the assumed regularity of $U$ and $V$, and recalling the l.s.c. of the entropy w.r.t. the weak convergence, the functional $\Phi$ is upper semicontinuous with respect to weak convergence. Therefore, the limit of such a subsequence is indeed a maximizer.
Let $(\overline{\mu},\overline{\nu})\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$ be a maximizer for $\Phi$. We note that $p\in\partial^{+}(U-\delta\mathcal{E})(\overline{\mu})$ has been already proved in [DS, Proposition 3.1].
The fact that $\xi$ belongs to the subdifferential $\partial^{-}V(\overline{\nu})$ follows directly from Proposition 2.14. Moreover, since $2\pi\delta\nabla|x|^{2}\in\mathrm{Tan}^{M}_{\overline{\nu}}\mathcal{P}_{2}(\mathbb{R}^{d})$, we also have $\xi\in\mathrm{Tan}^{K}_{\overline{\nu}}\mathcal{P}_{2}(\mathbb{R}^{d})$. Finally, observe that
$\|\gamma\|_{\overline{\nu}}\,\mathrm{Lip}(V;W_{2})\geq W_{2}(\overline{\nu},\exp_{\overline{\nu}}(\gamma))\,\mathrm{Lip}(V;W_{2})\geq V(\exp_{\overline{\nu}}(\gamma))-V(\overline{\nu})\geq\langle\xi,\gamma\rangle_{\theta,\overline{\nu}}+o(\|\gamma\|_{\overline{\nu}}),$
which yields the claim by choosing $\gamma=\frac{\varepsilon}{\|\xi\|_{\overline{\nu}}}\,\xi$ with $\varepsilon>0$, dividing by $\varepsilon$ and letting $\varepsilon\to 0$.
We now focus on the regularity of the optimizer $\overline{\mu}$. The proof is similar to that done in [DS, Propositon  3.7], and we just summarize the argument. Since $\mathrm{Lip}(U;W_{1})<\infty$, [DS, Lemma 3.3] yields $\overline{\mu}^{-1}\in L^{\infty}_{\text{loc}}.$ Therefore, $T_{\overline{\mu}}^{\overline{\nu}}$ is locally bounded, being a.e.-$\overline{\mu}$ the gradient of an everywhere defined convex function. In particular, $p$ is locally bounded. So that, we have
@@
$\begin{cases}p\in L_{\mathrm{loc}}^{\infty}(\mathbb{R}^{d})\\ p+\delta\nabla\log\overline{\mu}\in L^{\infty}(\mathbb{R}^{d})\end{cases}\Longrightarrow\nabla\log{\overline{\mu}}\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}^{d})\underbrace{\Longrightarrow}_{\overline{\mu},\overline{\mu}^{-1}\in L^{\infty}_{\mathrm{loc}}}\overline{\mu}\in W^{1,\infty}_{\mathrm{loc}}(\mathbb{R}^{d}).$
The following result will play a key role in obtaining a proper bound for the viscosity term.
@@
From the regularity assumptions on $\mu$, $\text{supp}(\mu)=\mathbb{R}^{d}.$. Therefore, it is sufficient to prove
@@
By Brenier Theorem, see [Bre, Proposition  3.1]) and the fact that $\text{supp}(\mu)=\mathbb{R}^{d}$, we have that $T_{\mu}^{\nu}=\nabla\phi$ $\mu$-a.e, where $\phi:\mathbb{R}^{d}\to\mathbb{R}$ is a finite convex function. Being $\phi$ locally Lipschitz, $T^{\nu}_{\mu}$ is locally bounded. We then regularize the transport map by a convolution kernel $\rho_{\alpha}:=\alpha^{-d}\rho(\frac{x}{\alpha})$, where $\rho\in C^{\infty}(\mathbb{R}^{d})$ is a probability density, compactly supported in the unitary ball $B_{1}(0)$ and $\alpha>0$. Set $T_{\alpha}:=\nabla\phi*\rho_{\alpha}=\nabla(\phi*\rho_{\alpha})$, then $T_{\alpha}$ inherits the local boundedness of $T_{\mu}^{\nu}$, and $T_{\alpha}$ converges to $T_{\mu}^{\nu}$ in $L^{2}_{\text{loc}}(\mathbb{R}^{d})$ (see [Evans, Theorem  6 pag. 630]). Since convolution preserves convexity, $\phi_{\alpha}:=\phi*\rho_{\alpha}$ is convex. Therefore, the following inequality holds for every $\alpha>0$
$\operatorname{div}(T_{\alpha}(x))=\text{Trace}(D^{2}\phi_{\alpha}(x))\geq 0\quad\forall x\in\mathbb{R}^{d}.$
@@
where $d(x,y)$ is the geodesic distance between $x$ and $y$ in $M$. See [petersen2006riemannian, Lemma  42]. In our context, the above inequality can be interpret as a weak analogue of
$\Delta_{\mu}\frac{1}{2}W_{2}^{2}(\mu,\nu):=\int_{\mathbb{R}^{d}}\Delta(\mathcal{D}_{\mu}\frac{1}{2}W_{2}^{2}(\mu,\nu)))d\mu\leq d.$
The inequality stated in Lemma 4.3 is a particular case of a more general result of independent interest, at least in the case $d\geq 2$.
@@
Suppose $d\geq 2$. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\mathcal{I}(\mu)<\infty$, $\mu^{-1}\in L^{\infty}_{\text{loc}}$ and $\mu\in W^{1,\infty}_{\text{loc}}$. Let $\mathcal{F}:\mathbb{R}^{d}\to\mathbb{R}$ be a $\Lambda$-geodesically concave functional and suppose $\mathcal{F}$ to be flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then
Recall that the squared Wasserstein distance is $1$-geodesically concave, i.e. the map $\mu\mapsto W_{2}^{2}(\mu,\nu)$ is 1-concave along Wasserstein geodesics for every $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Equivalently, $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a geodesic PC space, see [AGS, Definition 12.3.1] and later discussion. When $\mu$ is absolutely continuous, its flat derivative is given (up to an additive constant) by the optimal Kantorovich potential $\phi$ associated with the transport from $\mu$ to $\nu$. In particular, denoting by $T_{\mu}^{\nu}=\nabla\phi$ the optimal transport map, we have
@@
that explains why this proposition is a generalization in dimension $d\geq 2$ of Lemma 4.3.
Firstly, we derive a sufficient condition for a function on the Wasserstein space to have $\Lambda$-concave flat derivative, namely
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a flat differentiable function at a $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, s.t. there exists $\Lambda=\Lambda(\mu)$ for which the following inequality holds
@@
for all $y_{0},y_{1}\in\mathrm{supp}(\mu)$, with $y_{t}:=(1-t)y_{0}+ty_{1}\in\mathrm{supp}(\mu),\,t,h\in[0,1].$
Then $x\mapsto D_{\mu}\mathcal{F}(\mu,x)$ is $\Lambda$-convex in the classical $\mathbb{R}^{d}$ sense on $\mathrm{supp}(\mu)$.
@@
We derive the desired property from the explicit formula given in (2.12) for the flat derivative. Indeed, consider $y_{0},y_{1}\in\text{supp}(\mu)$ s.t. their convex interpolation $y_{t}=(1-t)y_{0}+ty_{1}\in\text{supp}(\mu)$ for all $t\in[0,1]$. Thanks to the assumption (4.4) on $\mathcal{F}$, and the representation formula (2.12), we have
@@
We now consider $\Lambda$-geodesically convex functions in the sense of the following definition.
Definition 4.7.
Given $\Lambda\in\mathbb{R}$, we say that $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ is $\Lambda$-geodesically convex w.r.t. $W_{2}$, if for any $\mu_{0},\mu_{1}\in\text{Dom}(\mathcal{F})$ and for every $W_{2}$ constant speed geodesic $t\mapsto\mu_{t}\in\text{Dom}(\mathcal{F})$ connecting $\mu_{0}$ and $\mu_{1}$, $\mathcal{F}$ is $\Lambda$-convex along this geodesic, i.e.
The following Lemma shows a geodesic convexity property along mixtures, i.e. flat convex combination of measures.
@@
  • In addition, if $\mathcal{F}$ is flat differentiable at $\mu$, then $\mathcal{F}$ satisfies the assumption (4.4) of Lemma 4.6.

### v2

@@ 4. Useful Estimates and Lemmata
We first start by showing that, as a consequence of the Benamou-Brenier control representation, the elements of the subdifferential of $W_{2}$-Lipschitz functions are bounded.
@@
($L^{2}$ version of [30, Lemma  3.4]) Let $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a $W_{2}$-Lipschitz function and set $L=\mathrm{Lip}(U;W_{2})<\infty$. Let $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
  • Let $\delta>0$, and $p\in\partial^{+}(U-\delta\mathcal{E})(\mu_{0})$ (resp. $p\in\partial^{-}(U+\delta\mathcal{E})(\mu_{0})$), then $\|p+\delta\nabla\log\mu_{0}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\leq L$ (resp. $\|p-\delta\nabla\log\mu_{0}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\leq L$)
  • Let $\delta=0$, and $\xi\in\partial^{+}U(\mu_{0})$ (resp. $\xi\in\partial^{-}U(\mu_{0})$), then $\|\xi\|_{\mu_{0}}\leq L$.
@@
We distinguish the two cases. The first case is treated very similarly to the one given in [30, Lemma  3.4], we only give a sketch of it in the superdifferential case.
@@
$\partial_{t}\mu_{t}+\operatorname{div}(v(x)\mu_{t})=0\quad\text{in }(t_{0},\infty),\quad\mu|_{t=t_{0}}=\mu_{0}.$
@@
Moreover, by the representation theorem for solutions of the continuity equation given in [2, Thm. 8.3.1] we have
@@
Moreover, by the renormalization property satisfied by $v$ and standard regularization and localization procedures (see [30, Lemma  3.4]), we also have
@@
Fix now $h,\varepsilon>0$ and $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$. By similar estimates performed in [30, Lemma  3.4], it follows, from the superdifferential property and the Cauchy Schwarz inequality, that for all sufficiently small $h>0$, depending on $\|v\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}$ and $\varepsilon$,
@@
is bounded in $L^{2}_{\mu_{0}}(\mathbb{R}^{d})$ with operator norm $L$. By the $L_{\mu_{0}}^{2}(\mathbb{R}^{d})$ duality, this element belongs to $L^{2}_{\mu_{0}}(\mathbb{R}^{d})$ and $\|p+\delta\nabla\log\mu_{0}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\leq L$.
In the case $\delta=0$, we argue as follows. If $\xi\in\partial^{+}U(\mu_{0})$, for $\gamma\in\mathcal{P}_{2,\overline{\mu}_{0}}(\mathbb{R}^{2d})$ and $\theta\in\Gamma_{\mu_{0}}(\xi,\gamma)$,
$\|\gamma\|_{\mu_{0}}\,\mathrm{Lip}(U;W_{2})\geq W_{2}(\mu_{0},\exp_{\mu_{0}}(\gamma))\,\mathrm{Lip}(U;W_{2})\geq U(\exp_{\mu_{0}}(\gamma))-U(\mu_{0})\geq\langle\xi,\gamma\rangle_{\theta,\mu_{0}}+o(\|\gamma\|_{\mu_{0}}),$
which yields the claim by choosing $\gamma=\frac{\varepsilon}{\|\xi\|_{\mu_{0}}}\,\xi$ with $\varepsilon>0$, dividing by $\varepsilon$ and letting $\varepsilon\to 0$.
Viscosity solutions of (1.1) enjoy a time regularization.
Lemma 4.2.
Let $U$ be a viscosity solution of (1.1). Then the following implication holds
$L:=\sup_{t\in[0,T]}\mathrm{Lip}(U(t,\cdot);W_{2})\Longrightarrow\sup_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\mathrm{Lip}_{t}(U(\cdot,\mu))\leq\sup_{\{\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d}):\|\gamma\|{{}_{(\pi_{1})_{\#}\gamma}}\leq L\}}|\mathcal{H}(\gamma)|<\infty.$
Proof.
It is sufficient to prove that each element in the time-differential is bounded by the constant $\sup_{\{\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d})\,:\,\|\gamma\|_{(\pi_{1})_{\#}\gamma}\leq L\}}|\mathcal{H}(\gamma)|$. Firstly, we prove that this constant is finite. Indeed, by the growth assumption on $H$, using (2.20), we have
$\sup_{\{\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d}):\|\gamma\|_{(\pi_{1})_{\#}\gamma}\leq L\}}|\mathcal{H}(\gamma)|\leq\sup_{\{\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d}):\|\gamma\|_{(\pi_{1})_{\#}\gamma}\leq L\}}C(1+\|\gamma\|^{2}_{{(\pi_{1})_{\#}\gamma}})\leq C(1+L^{2}).$
Now fix $(t,\mu)\in[0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and consider $(r,\gamma)\in\partial^{+}U(t,\mu)$. By Definition 2.19, being $U$ a viscosity subsolution of (1.1), we have
[…]
Since $\gamma$ is an element of the superdifferential of a uniform in time Lipschitz function, we have
$\|\gamma\|_{\mu}\leq L.$
Therefore, taking the infimum on a bigger set, we have
$\inf_{\{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}):\|\gamma\|_{\mu}\leq L\}}-|\mathcal{H}(\gamma)|\leq r\quad\forall r\in\partial_{t}^{+}U(t,\mu).$
With a similar computation, using the supersolution property, one can prove an analogous upper bound for the elements in the subdifferential. Namely,
$r\leq\sup_{\{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}):\|\gamma\|_{\mu}\leq L\}}|\mathcal{H}(\gamma)|,\quad\forall r\in\partial_{t}^{-}U(t,\mu).$
We now state a result that will be used in the proof of the vanishing viscosity limit.
Lemma 4.3 (Analogue of Proposition 3.1 of [30] in our context).
Let $U,V:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be two bounded functions s.t. $\mathrm{Lip}(U;W_{1})+\mathrm{Lip}(V;W_{2})<\infty$ and suppose $V$ is weakly l.s.c.. Fix $\alpha,\delta>0$, then the function $\Phi:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})\to[-\infty,+\infty)$ defined by
$\Phi(\mu,\nu):=U(\mu)-V(\nu)-\frac{1}{2\alpha}W_{2}(\mu,\nu)-\delta\mathcal{E}(\mu)-\delta\pi\mathcal{M}_{2}(\mu)-\delta\pi\mathcal{M}_{2}(\nu)$
@@
    where $\sigma\in\Gamma_{0}(\overline{\nu},\overline{\mu})$ is such that $(\pi_{2},\pi_{1})_{\#}\sigma=(\mathrm{Id}\times T_{\overline{\mu}}^{\overline{\nu}})_{\#}\overline{\mu}\in\Gamma_{0}(\overline{\mu},\overline{\nu}),$ then $p\in{\partial^{+}}\big(U-\delta\mathcal{E}\big)(\overline{\mu})$ and $\xi\in\partial^{-}V(\overline{\nu})$. Moreover,
    $\|p+\delta\nabla\log\overline{\mu}\|_{L^{\infty}(\mathbb{R}^{d})}\leq\mathrm{Lip}(U;W_{1}),\,\|p+\delta\nabla\log\overline{\mu}\|_{L^{2}_{\overline{\mu}}(\mathbb{R}^{d})}\leq\mathrm{Lip}(U;W_{2})\text{ and }\|\xi\|_{\overline{\nu}}\leq\mathrm{Lip}(V;W_{2});$
@@
The attainment of the maximum follows the same argument used in the proof of [30, Proposition 3.1]. The moment penalization ensures that any maximizing sequence $(\mu_{n},\nu_{n})$ remains in a compact subset of $\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$ in the weak topology (2.17), more precisely $\sup_{n}\mathcal{M}_{2}(\mu_{n})+\mathcal{M}_{2}(\nu_{n})<\infty$. As a consequence, there exists a (not relabeled) subsequence $(\mu_{n},\nu_{n})$ converging weakly to some $(\overline{\mu},\overline{\nu})\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$. By the assumed regularity of $U$ and $V$, together with the lower semicontinuity of the entropy and moments with respect to narrow convergence, the functional $\Phi$ is upper semicontinuous with respect to the weak topology. Consequently, the limit pair $(\overline{\mu},\overline{\nu})$ is indeed a maximizer of $\Phi$.
Let $(\overline{\mu},\overline{\nu})\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$ be a maximizer for $\Phi$. We note that $p\in\partial^{+}(U-\delta\mathcal{E})(\overline{\mu})$ has been already proved in [30, Proposition 3.1].
The fact that $\xi$ belongs to the subdifferential $\partial^{-}V(\overline{\nu})$ follows directly from Proposition 2.15. Moreover, since $2\pi\delta\nabla|x|^{2}\in\mathrm{Tan}^{M}_{\overline{\nu}}\mathcal{P}_{2}(\mathbb{R}^{d})$, we also have $\xi\in\mathrm{Tan}^{K}_{\overline{\nu}}\mathcal{P}_{2}(\mathbb{R}^{d})$. The bound $\|\xi\|_{\overline{\nu}}\leq\mathrm{Lip}(V;W_{2})$ is a consequence of Lemma 4.1.
We now focus on the regularity of the optimizer $\overline{\mu}$. The proof is similar to that done in [30, Proposition  3.1], and we just summarize the argument. Since $\mathrm{Lip}(U;W_{1})<\infty$, [30, Lemma 3.3] yields $\overline{\mu}^{-1}\in L^{\infty}_{\text{loc}}(\mathbb{R}^{d}).$ Therefore, $T_{\overline{\mu}}^{\overline{\nu}}$ is locally bounded, being a.e.-$\overline{\mu}$ the gradient of an everywhere defined convex function. In particular, $p$ is locally bounded. So that, we have
@@
$\begin{cases}p\in L_{\mathrm{loc}}^{\infty}(\mathbb{R}^{d})\\ p+\delta\nabla\log\overline{\mu}\in L^{\infty}(\mathbb{R}^{d})\end{cases}\Longrightarrow\nabla\log{\overline{\mu}}\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}^{d})\underbrace{\Longrightarrow}_{\overline{\mu},\overline{\mu}^{-1}\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}^{d})}\overline{\mu},\overline{\mu}^{-1}\in W^{1,\infty}_{\mathrm{loc}}(\mathbb{R}^{d}).$
The following result will play a key role in obtaining a proper bound for the viscosity term in the proof of the vanishing viscosity limit.
@@
From the regularity assumptions on $\mu$, it follows $\text{supp}(\mu)=\mathbb{R}^{d}$. Therefore, it is sufficient to prove
@@
By Brenier Theorem, see [10, Proposition  3.1], and the fact that $\text{supp}(\mu)=\mathbb{R}^{d}$, we have that $T_{\mu}^{\nu}=\nabla\phi$ $\mu$-a.e, where $\phi:\mathbb{R}^{d}\to\mathbb{R}$ is a finite convex function. Being $\phi$ locally Lipschitz, $T^{\nu}_{\mu}$ is locally bounded. We then regularize the transport map by a convolution kernel $\rho_{\alpha}:=\alpha^{-d}\rho(\frac{x}{\alpha})$, where $\rho\in C^{\infty}(\mathbb{R}^{d})$ is a probability density, compactly supported in the unitary ball $B_{1}(0)$ and $\alpha>0$. Set $T_{\alpha}:=\nabla\phi*\rho_{\alpha}=\nabla(\phi*\rho_{\alpha})$, then $T_{\alpha}$ inherits the local boundedness of $T_{\mu}^{\nu}$, and $T_{\alpha}$ converges to $T_{\mu}^{\nu}$ in $L^{2}_{\text{loc}}(\mathbb{R}^{d})$ (see [32, Theorem  6 pag. 630]). Since convolution preserves convexity, $\phi_{\alpha}:=\phi*\rho_{\alpha}$ is convex. Therefore, the following inequality holds for every $\alpha>0$
$\operatorname{div}(T_{\alpha}(x))=\text{tr}(D^{2}\phi_{\alpha}(x))\geq 0\quad\forall x\in\mathbb{R}^{d}.$
@@
where $d(x,y)$ is the geodesic distance between $x$ and $y$ in $M$. See [44, Lemma  42]. In our context, the above inequality can be interpret as a weak analogue of
$\Delta_{\mu}\frac{1}{2}W_{2}^{2}(\mu,\nu):=\int_{\mathbb{R}^{d}}\Delta_{x}(\mathcal{D}_{\mu}\frac{1}{2}W_{2}^{2}(\cdot,\nu)(\mu,x))d\mu(x)\leq d.$
The inequality stated in Lemma 4.4 is a particular case of a more general result of independent interest, that holds at least when $d\geq 2$ for $\Lambda$-geodesically concave functions. Before stating this result, let us recall the following definition.
Definition 4.6.
Given $\Lambda\in\mathbb{R}$, we say that $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ is $\Lambda$-geodesically convex w.r.t. $W_{2}$, if for any $\mu_{0},\mu_{1}\in\text{Dom}(\mathcal{F})$ and for every $W_{2}$-constant speed geodesic $t\mapsto\mu_{t}\in\text{Dom}(\mathcal{F})$ connecting $\mu_{0}$ and $\mu_{1}$, $\mathcal{F}$ is $\Lambda$-convex along this geodesic, i.e.
@@
Suppose $d\geq 2$. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\mathcal{I}(\mu)<\infty$, $\mu^{-1}\in L^{\infty}_{\text{loc}}(\mathbb{R}^{d})$ and $\mu\in W^{1,\infty}_{\text{loc}}(\mathbb{R}^{d})$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a $\Lambda$-geodesically concave functional and suppose $\mathcal{F}$ to be flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then $\mathbb{R}^{d}\ni x\mapsto\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$ is $\Lambda$-concave, in particular $\mathbb{R}^{d}\ni x\mapsto\nabla\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$ exists $\mu$-a.e.. Moreover
$\int_{\mathbb{R}^{d}}\big<\nabla\log\mu(x),\nabla\mathcal{D}_{\mu}\mathcal{F}(\mu,x)\big>d\mu(x)\leq d\Lambda.$
Before proving the above result, let us introduce the context. Recall that the squared Wasserstein distance is $1$-geodesically concave, i.e. the map $\mu\mapsto W_{2}^{2}(\mu,\nu)$ is 1-concave along Wasserstein geodesics for every $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Equivalently, $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a geodesic PC space, see [2, Definition 12.3.1] and later discussions. When $\mu$ is absolutely continuous, the flat derivative of $\mu\mapsto W_{2}^{2}(\mu,\nu)$ is given (up to an additive constant) by the optimal Kantorovich potential $\phi$ associated with the transport from $\mu$ to $\nu$. In particular, denoting by $T_{\mu}^{\nu}=\nabla\phi$ the optimal transport map, we have
@@
that explains why the above proposition is a generalization in dimension $d\geq 2$ of Lemma 4.4.
Firstly, we derive, for functions on the Wasserstein space, a sufficient condition to have a $\Lambda$-convex flat derivative, namely
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a flat differentiable function at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, s.t. there exists $\Lambda=\Lambda(\mu)$ for which the following inequality holds
@@
for all $y_{0},y_{1}\in\mathbb{R}^{d}$, with $y_{t}:=(1-t)y_{0}+ty_{1},\,t,h\in[0,1].$
Then, $\mathbb{R}^{d}\ni x\mapsto\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$ is $\Lambda$-convex.
@@
We derive the desired property from the explicit formula given in (2.14) for the flat derivative. Indeed, consider $y_{0},y_{1}$ and their convex interpolation $y_{t}=(1-t)y_{0}+ty_{1}$ for all $t\in[0,1]$. Thanks to the assumption (4.3) on $\mathcal{F}$, and the representation formula (2.14), we have
@@
Secondly, the following Lemma shows a geodesic convexity property along mixtures, i.e. flat convex combination of measures.
@@
  • (2)
    In addition, if $\mathcal{F}$ is flat differentiable at $\mu$, then Lemma 4.8 holds.